# General Riemann Integration on Partially Ordered Net Convergence Structures

### v1

@@ 3 Partially Ordered Net Convergence Structures and Riemann Integration
We first fix, following the terminology and definitions of [27], the space $\mathcal{N}(X)$ as the set of all admissible nets (see the next comments) in an abstract space $X$ which will always denote, except when specified, a Riesz space - even though most results of this section, and work, are valid for lattice ordered groups or semigroups. It is important to point out that the concept of an admissible net is of a technical nature (based on the axioms of ZFC, as explained in [27]), which we are not going to emphasize much - it is always possible to find, for each net, an admissible net such that the two have equal limit in the sense of the next definition. So, we shall work only with this class (i.e, identify nets with its admissible class) and comment no further.
To give the essential definition, we need (see Page 4 of [27]):
@@
Then, we have the following main definition from [27]. As in further constructions, $I$ will, in general, denote a general directed set that is to serve as the domain of the nets.
@@
which is defined by the relation $\{x_{i}\}_{i\in I}\in\eta(x)$, written sometimes as the limit notation $\{x_{i}\}_{i\in I}\rightarrow x$, $\lim_{(I,\eta)}x_{i}=x$ or $\lim_{i}x_{i}=x$, such that:
@@
    If a net $\{x_{i}\}_{i\in I}$ is constant and equals to a $x\in X$, then $\{x_{i}\}_{i\in I}\in\eta(x)$ and its limit is $a$;
@@
For further comparison with other convergences, the following result of [27] is of interest:
@@
    If $\{x_{i}\}_{i\in I}$ converges to $x$ and $\{y_{j}\}_{j\in J}$ is tail equivalent to $\{x_{i}\}_{i\in I}$, then $\{y_{j}\}_{j\in J}$ converges to $x$.
Thus, this result shows that this general convergence concept includes the net convergence definition of [28]. It can also be seen that it includes the net convergences of [12], based on axioms of pages 306-307 from it, and [40].
The integral that will be defined below needs only the three axioms of Definition 3.2 of a net convergence structure, but will be supplemented by the following conditions from [12], first appearing in the work of Boccuto and Candeloro [57], which will be used for uniformity with the approaches of the aforementioned authors and integration in partially ordered spaces more generally:
Definition 3.3 (Boccuto-Candeloro Net Convergence Structure).
A Bocutto-Candeloro net convergence structure on (a Riesz space) $X$ is a function
$\eta:X\rightarrow 2^{\mathcal{N}(X)},$
which is defined by the relation $\{x_{i}\}_{i\in I}\in\eta(x)$, written sometimes as the limit notation $\{x_{i}\}_{i\in I}\rightarrow x$, $\lim_{(I,\eta)}x_{i}=x$ or $\lim_{i}x_{i}=x$, such that:
  1. (i)
    If $\{x_{i}\}_{i\in I}\rightarrow x$ and $\{y_{j}\}_{j\in j}\rightarrow y$, then $\{x_{i}+y_{j}\}_{(i,j)\in I\times J}\rightarrow x+y$; if $\{x_{i}\}_{i\in I}\rightarrow x$ and $\{y_{i}\}_{i\in I}\rightarrow y$ are nets such that $x_{i}\leq y_{i}$ for all $i\in I$, or in a cofinal subset of $I$, then $x\leq y$.
  2. (ii)
    If $\{x_{i}\}_{i\in I}\rightarrow x$, then $\{|x_{i}|\}_{i\in I}\rightarrow|x|$.
  3. (iii)
    If $\{x_{i}\}_{i\in I}\rightarrow x$ and $\{y_{i}\}_{i\in I}\rightarrow x$ and $\{z_{i}\}_{i\in I}$ is a net such that there exists $i_{0}\in I$ such that for every $i\geq i_{0}$, $x_{i}\leq z_{i}\leq y_{i}$, then $\{z_{i}\}_{i\in I}\rightarrow x$.
  4. (iv)
    If $\{x_{n}\}_{i\in\mathbb{N}}\rightarrow x$ is a (convergent) sequence, then $\{\frac{1}{n}x_{n}\}_{n\in\mathbb{N}}\rightarrow 0$
  5. (v)
    If a net $\{x_{i}\}_{i\in I}$ is constant and equals to a $x\in X$, then $\{x_{i}\}_{i\in I}\rightarrow a$.
  6. (vi)
    If a net converges to $x$, every quasi-subnet of it converges to $x$.
  7. (vii)
We will also assume, unless stated to the contrary, that all our convergence structures on the Riesz space $X$ will be Hausdorff: that is, each net admits, at most, only one limit. We refer to [27] and [26] for further details and other interesting restrictions.
We will also make use of the following useful definition of a Cauchy type partially ordered convergence (see, for example, [26]):
@@ A Bocutto-Candeloro net convergence structure on (a Riesz space) $X$ is a function
Let $(X,\eta)$ be a Riesz space equipped with a Boccuto-Candeloro net convergent structure $\eta$. We say that a net $\{x_{i}\}_{i\in I}$ is Cauchy if its associated difference
@@
is convergent to 0 in the Boccuto-Candeloro net convergence structure $\eta$.
If each Cauchy net in $X$ associated to $\eta$ is convergent [^2] Notice that every convergent sequence is Cauchy in an arbitrary net convergence structure. For details, see [27]., the convergence structure $\eta$ is called Cauchy or, equivalently, $X$ is called complete (in reference to the net convergence structure $\eta$).
Concerning the notions of convergence in partially ordered spaces defined in the previous section, we have, by a consequence of Definition 9.1 and Proposition 9.2 of [27]:
Proposition 3.2 (Equivalence of Convergences in Riesz spaces).
Let $X$ be a partially ordered space. Then, the $(o_{3})$, $(o)$ and $(D)$-convergences are Boccuto-Candeloro net convergence structures. If $X$ is, further, a lattice ordered group, the first two are equivalent.
Further, we have the following result (see [11]):
@@
In a lattice ordered group or Riesz space $X$, order convergence is Cauchy.
@@
In a lattice ordered group or Riesz space $X$ that is weakly $\sigma$-distributive and Dedekind complete, (D)-convergence is Cauchy.
@@
As we have not found in the literature a proof for this result for general nets, we supply a proof inspired in the result for integrals in [58].
@@
$|x_{i_{1}}-x_{i_{2}}|\leq\bigvee_{n=1}^{\infty}a_{i\varphi(i)},$
@@
As a consequence, as for every $i\geq i^{*}(\varphi)$,
@@
3.1 The Net Riemann Integral
Now, to begin the preparation for integration structures in the context of abstract net convergence of the type above, we need a further notion from the general convergence theory (see [27] and [26]):
Definition 3.5 (Continuous Maps Between Convergent Spaces).
Let $X$ and $Y$ be Riesz spaces with net convergence structures $\eta_{1}$ and $\eta_{2}$ respectively. We say that a map $f:X\rightarrow Y$ is continuous if:
for every net convergent net $\{x_{i}\}_{i\in I}\in\mathcal{N}(X)$.
To integrate measure and functions, we will need, in conjunction to a Boccuto-Candeloro net convergence structures in each image space of these objects, a bilinear structure that connects the two, one on which the measure lives, the other where the functions live, to a third space equipped with a convergence structure of the cited type, with the additional requirement that it preserves order (isotonicity) and continuity of the algebraic operations. More precisely, we suppose the existence of a bilinear "product" as in [30] (see also [38], [11], [59] and [31]):
Definition 3.6 (Product in Riesz spaces).
let $(X,\eta_{1})$, $(Y,\eta_{2})$ and $(Z,\eta_{3})$ be three Riesz spaces with respective Boccuto-Candeloro convergence structures $\eta_{i}$, $i=1,2,3$.
@@ 3.1 The Net Riemann Integral
    Isotone on the positive cones of $X$ and $Y$, $X^{+}$ and $Y^{+}$ respectively,
@@
    are continuous in the associated Boccuto-Candeloro net convergence structures of the spaces in each case,
is called a bilinear product, or only product.
For the measure-theoretic structures which the integration theory in the present work will be reduced to, and were constructed by [21], [22], [23], [24] in the (Banach or uniform semigroup) topological space-valued case of measures and functions (and products), we use the next construction:
@@
Let $(\Omega,\mathcal{H})$ a fixed paved space. A net:
@@
Given these two definitions, we have the main structural object of the present section, defined in analogy to Definition 1.5 of Szabó and Száz [22]:
Definition 3.8 (Partially Ordered Integration Structure).
@@
    $(X,\eta_{1}),(Y,\eta_{2}),(Z,\eta_{3})$, Riesz spaces with Boccuto-Candeloro convergence structures connected by a product $\bullet$;
will be called an partially ordered integration system.
With these definitions, we may state the main object of our integration theory, again borrowing ideas of Szabó and Száz, and Fleischer:
@@
Let $f:\Omega\rightarrow X$ be a function and $\mu:\mathcal{H}\rightarrow Y$ be a set function, which will be called a measure, both given with respect to a partially ordered integration structure
@@
Then, defining as:
$S_{\gamma}(f,\mu)=\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}}),$
@@
an abstract Riemann sum in $Z$, if the following limit exists,
$\int_{\Omega}fd\mu=\lim_{(\Gamma,\eta_{3})}S_{\gamma}(f,\mu),$
@@
it will be called the Net Riemann integral of $f$ with respect to $\mu$ on $\Omega$, or only the Net Riemann integral of $f$ if the set function is uniquely specified.
For a set $A\in\mathcal{H}$, we say that $f$ is net Riemann integrable on $A$ if:
$\int_{\Omega}f\mathbbm{1}_{A}d\mu,$
exists and in this case define:
$\int_{A}f\mu=\int_{\Omega}f\mathbbm{1}_{A}d\mu$
and note that, in this case, the Riemann sums cannot, in general, be restricted to $\sigma$ such that all its elements are contained in $A$.
We now comment briefly on a possible modification of this integral inspired by [24].
Observation 3.1 (Fleischer [24] Integration Domains and Fleischer Net Riemann Integral).
In general, the integration process above doesn’t take into account specific properties of the set $A\in\mathcal{H}$ in integrating a function $f$ on $A$, in the sense that the Fleischer-Szabó-Száz net of integration is independent of the set $A$. We can refine the integration concept above and ask that, for each $A\in\mathcal{H}$, we have a net in the space of finite collection of sets and points as specified above, defined on a directed set $\Gamma_{A}$. Thus, we obtain a collection of directed sets $\{\Gamma_{A}\}_{A\in\mathcal{H}}$ and Fleischer-Szabó-Száz nets associated with them. The integral of $f$ on $A$ is then the limit along $\Gamma_{A}$ of the Riemann sums of $f$ associated to Fleischer-Szabó-Száz net of $\Gamma_{A}$. This notion of integral will be called the Fleishcer Net Riemann integral in honor of Fleischer [24] contribution and to distinguish it from the construction above.
Questions of the additivity of the set function induced by such integration procedure, as well as subsets integration (i.e, existence on the integral of $f$ on $B$ implies existence on $A$ for each $A\subseteq B$ and $A,B\in\mathcal{H}$), becomes more delicate to answer than for the integral of Definition 3.9. Nevertheless, most basic properties, such as additivity, isotonocity, and others, remain the same for this integral. Therefore, we focus on the basic Net Riemann integral, and offer some comments on Fleischer’s one when appropriate, specially in the last sections.
@@
Following Definition 3.9, we denote by $L^{1}(\mu)$ the space of all functions $f:\Omega\rightarrow X$ Net Riemann integrable on $\Omega$  [^3] Most of the properties below are valid for $f$ integrable on $A\in\mathcal{H}$ with suitable modifications. In general, we also omit the ”on $\Omega$” if it causes no confusion. with respect to the (fixed) set function $\mu:\Omega\rightarrow Y$ in the context of a (fixed) partially ordered integration structure.
Then, with the previous structures in mind, before presenting examples of specific integration procedures that fit in this archetype, we collect general properties of the Net Riemann integral that can be gathered essentially from the definition and without imposing too many conditions on $f$ and/or $\mu$ and its associated spaces:
Theorem 3.1 (Properties of the Net Riemann integral).
The Net Riemann integral has the following properties:
@@
    Let $f,g\in L^{1}(\mu)$. Then,
    $\int_{\Omega}(f+g)d\mu=\int_{\Omega}fd\mu+\int_{\Omega}gd\mu,$
    i.e, the integral is additive on the space $L^{1}(\mu)$;
@@
    Let $f\in L^{1}(\mu)$ such that $f\geq 0$ and $\mu\geq 0$. Then,
    $\int_{\Omega}fd\mu\geq 0,$
  3. (iii)
    Let $\mu\geq 0$ and $f,g\in L^{1}(\mu)$ such that:
@@
    $\int_{\Omega}gd\mu\geq\int_{\Omega}fd\mu,$
    i.e, the integral is an isotone map on $L^{1}(\mu)$ equipped with the partial ordering inherited from $X$.
  4. (iv)
    Let $\mu\geq 0$ and $f\in L^{1}(\mu)$ such that $|f|\in L^{1}(\mu)$, then:
    $|\int_{\Omega}fd\mu|\leq\int_{\Omega}|f|d\mu.$
@@
(i) Let
$S_{\gamma}(f,\mu),$
$S_{\gamma}(g,\mu),$
be the Riemann sums of $f$ and $g$ corresponding to the (fixed) integration structure of $L^{1}(\mu)$. Then, it is readily seen that, as the product is bilinear:
$S(f+g,\mu)=\sum_{i\in I_{\gamma}}(f+g)(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})=\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})+\sum_{i\in I_{\gamma}}g(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}})=S(f,\mu)+S(g,\mu),$
which, by taking the limit with respect to the Boccuto-Candeloro net convergence structure $\eta_{3}$, and using axiom $(i)$ of it, it follows that:
$\lim_{(\Gamma,\eta_{3})}S_{\gamma}((f+g),\mu)=\lim_{(\Gamma,\eta_{3})}S_{\gamma}(f,\mu)+\lim_{(\Gamma,\eta_{3})}S_{\gamma}(g,\mu),$
which is the same as:
$\int_{\Omega}(f+g)d\mu=\int_{\Omega}fd\mu+\int_{\Omega}gd\mu$
(ii) Let $f\in L^{1}$ such that it, and $\mu$, are non-negative (in $X$ and $Y$ respectively). Let the Riemann sum of $f$ in the fixed integration structure given by $S_{\gamma}(f,\mu)$. By the isotone property of the product and the non-negativity of both $f$ and $\mu$, it follows that:
@@
i.e, the sum is non-negative in $Z$ for every $I_{\gamma}$ and $\gamma\in\Gamma$. Therefore, by taking the limit with respect to the Boccuto-Candeloro net convergence structure $\eta_{3}$, and using the second component of axiom $(i)$, and using axiom $(v)$, it follows that:

### v2

@@ Abstract
In this sense, following O’Brien et al. (2023), we fix the space $\mathcal{N}(X)$ as the set of all admissible nets in an abstract (non-empty) space $X$. The concept of an admissible net is of a technical nature (based on the axioms of ZFC, as explained in O’Brien et al. (2023)), which we are not going to emphasize much - it is always possible to find, for each net, an admissible net such that the two have equal limit in the sense of the next definition. So, we shall work only with this class (i.e, identify nets with its admissible class) and comment no further.
In the rest of this work $I$ and $L$ will, in general, denote general directed (upwards or downwards) sets that are to serve as the domain of the nets. We shall not refrain to use other (Latin) indexes when these two are already in use in the same context.
We begin with the following preliminary concept (see Page 4 of O’Brien et al. (2023)):
@@
Then, we have the following main definition from O’Brien et al. (2023).
@@
which is defined by the relation $\{x_{i}\}_{i\in I}\in\eta(x)$, written sometimes as the limit notation $\{x_{i}\}_{i\in I}\rightarrow x$, $\lim_{(I,\eta)}x_{i}=x$ or $\lim_{i}x_{i}=x$, such that $x$ is called the limit of the net, and:
@@
    If a net $\{x_{i}\}_{i\in I}$ is constant and equals to a $x\in X$, then $\{x_{i}\}_{i\in I}\in\eta(x)$ and its limit is $x$;
@@
We call $(X,\eta)$, or only $X$ if the convergence is implicit, a net convergence space.
If the space in question has a partial order, we also have the following analogous construction of Bilokopytov et al. (2024), which is a generalization of the various order convergences above, as well as of relative uniform, $(uo),(un)$ and $(uaw)$ convergences:
Definition 1.14 (Locally Solid or Partially Ordered Net Convergence Structure).
Given a net convergence space $(X,\eta)$ where $X$ is a Riesz space, we say that $\eta$ is locally solid if, besides the axioms of Definition 1.13, it also satisfies:
  1. 1.
    If $\{y_{i}\}_{i\in I}$ converges to $y$ in $X$ and $\{x_{i}\}_{i\in I}$ is a net such that:
    $|x_{i}|\leq|y_{i}|,$
    for every $i\in I$, then $\{x_{i}\}_{i\in I}$ converges to $y$.
  2. 2.
    If $\{y_{l}\}_{l\in L}$ converges to $y$ in $X$ and $\{x_{i}\}_{i\in I}$ is a net such that for every $l_{0}\in I$ there exists $i_{0}\in L$ such that for every $i\geq i_{0}$ there exists $l\geq l_{0}$ with $|x_{i}|\leq|y_{l}|$, then $\{x_{i}\}_{i\in I}$ converges to $y$.
We shall also say that $(X,\eta)$ is, in this case, a partially ordered net convergence structure.
We now have an analogous definition of the Hausdorff separation property:
Definition 1.15 (Hausdorff Net Convergence Spaces).
We say that a net convergence space $(X,\eta)$ is Hausdorff if each net has only one limit.
We now add the following hypothesis for the rest of this work:
Hypothesis 1.1 (Hausdorff Hypothesis).
From now, all the net convergence spaces in this work shall be supposed, or proved, to be Hausdorff.
For further comparison with other convergences, the following result of O’Brien et al. (2023) is of interest:
@@
    If $\{x_{i}\}_{i\in I}$ converges to $x$ and $\{y_{i^{{}^{\prime}}}\}_{i^{{}^{\prime}}\in I^{{}^{\prime}}}$ is tail equivalent to $\{x_{i}\}_{i\in I}$, then $\{y_{i^{{}^{\prime}}}\}_{i^{{}^{\prime}}\in I^{{}^{\prime}}}$ converges to $x$.
Thus, this result shows that this general convergence concept includes the net convergence definition of Higgs (1988). It can also be seen that it includes the net convergences of Boccuto and Dimitriou (2015), based on axioms of pages 306-307 from it, and Boccuto and Candeloro (2009a), as well as the other convergences cited above, which also have filter analogues.
@@
Let $(X,\eta)$ be a net convergence space. We say that a net $\{x_{i}\}_{i\in I}$ is Cauchy if its associated difference
@@
is convergent to 0 in the net convergence structure $\eta$.
If each Cauchy net in $X$ associated to $\eta$ is convergent [^4] Notice that every convergent sequence is Cauchy in an arbitrary net convergence structure. For details, see O’Brien et al. (2023)., the convergence structure $\eta$ is called Cauchy or, equivalently, $X$ is called complete (in reference to the net convergence structure $\eta$).
Continuity is an important concept, and it can be generalized to this case (see O’Brien et al. (2023) and Bilokopytov et al. (2024)):
Definition 1.17 (Continuous Maps Between Convergent Spaces).
Let $X$ and $Y$ be net convergence spaces with net convergence structures $\eta_{1}$ and $\eta_{2}$ respectively. We say that a map $f:X\rightarrow Y$ is continuous if and only if:
for every net convergent net $\{x_{i}\}_{i\in I}$ in $X$.
With regards to this last definition, we shall add to the above the following two hypothesis for the rest of this work:
Hypothesis 1.2 (Hypothesis and Definition - Continuity of Addition for Vector Spaces and Product Convergence Structure).
Let $(X,\eta)$ be a net convergence space with $X$ being a vector space. For the rest of this work, we shall suppose that the operation of addition and multiplication by real numbers for each fixed $r\in\mathbb{R}$ in $X$ are continuous. That is:
[…]
are both continuous, where $X\times X$ is equipped with what we call the product convergence structure, that is: a net $\{x_{i}\}_{i\in I}$ of the form $x_{i}=(x_{i}^{1},x_{i}^{2})$ is convergence to $x=(x_{1},x_{2})$ in $X$ if and only if the marginal nets $\{x_{i}^{1}\}_{i\in I}$ and $\{x_{i}^{2}\}_{i\in I}$ converge to $x_{1}$ and $x_{2}$ in $X$ with respect to $\eta$ respectively. The product convergence structure in an infinite (possibly uncountable) product is defined in the analogous (i.e, convergence in each projection) way, with addition and multiplication by (real) constants understood in the coordinate-to-coordinate sense.
This is related to the concept of linear convergence structure as in O’Brien et al. (2023).
Hypothesis 1.3 (Hypothesis - Locally Solid Convergence).
Let $(X,\eta)$ be a net convergence space with $X$ being a Riesz space. For the rest of this work, we shall suppose that the the convergence structure $\eta$ in $X$ is locally solid.
To connect these concepts with the ones in the previous section, we note the next result due to McGill (1975):
@@
In a Dedekind complete Riesz space $X$, $(o)$-convergence is Cauchy.
@@
In a Riesz space $X$ that is weakly $\sigma$-distributive and Dedekind complete, (D)-convergence is Cauchy.
@@
As we have not found in the literature a proof for this result for general nets, we supply a proof inspired in the result for integrals in Riečan and Vrábelová (1993).
@@
$|x_{i_{1}}-x_{i_{2}}|\leq\bigvee_{n=1}^{\infty}a_{n\varphi(n)},$
@@
As a consequence, for every $i\geq i^{*}(\varphi)$,
@@
This gives all we need to define the main integration procedure of this work, which we now do.
2 The Abstract Net Riemann Integral.
The convergence ideas above can give rise to a very general integration procedure when combined with an algebraic element of some sort. Basically, this combination of an algebraic operation and a convergence concept encapsulates the notion that defining a Riemann type integral is a matter of having a notion of well defined sums (or products) and an approximation procedure. That is, fixing $(X,\eta,\star)$ a net convergence structure with an operation $\star$ compatible with the convergence (in some sense), we may consider integration (of a function against a measure) as a limit of a net/filter of formal sums of the form:
$(f,\mu)\rightarrow\sum_{i\in I_{\gamma}}f(\tau)\mu(\sigma),$
where $f$ and $\mu$ have values in spaces that can be connected with $X$ by some map (not necessarily linear).
We now follow this intuition to define a very general integral that has some interesting properties, and which was considered, in a less general setting and with a narrower definition in Szabó and Száz (1988) and Fleischer (1990). After defining our general integration procedure, we shown that most of the integrals in the literature are special cases of ours, and we also generalize specific convergence/additivity theorems of these integrals to the present abstraction. We present most of the results in the language of nets, but the dual filter notation is also be useful.
As we will be dealing with different types of integrals in this work, we identify them by an superscript consisting of an upper case letter with a direct summary of the name, e.g. $\int^{R}$ for the Riemann integral. The only exception will be the main integral of this section, which will be denoted without superscript.
In general, we will be integrating measures and functions with values in different spaces in conjunction to net convergence structures in each of these. Therefore, a concatenation structure that connects the two to a third space equipped with a convergence structure of the cited type, with the additional requirement of continuity of the algebraic operations, is needed. More precisely, we suppose the existence of a "bilinear product" as in Pavlakos (1983), Dinculeanu (2014), Dinculeanu (2000) and other classics of vector integration (see also Boccuto (1995), Boccuto et al. (2012b), Filter (1994) and Pavlakos (1991)):
Definition 2.1 (Bilinear Product in Net Convergence Spaces).
let $(X,\eta_{1})$, $(Y,\eta_{2})$ and $(Z,\eta_{3})$ be three vector spaces with respective net convergence structures $\eta_{i}$, $i=1,2,3$.
@@
    Bilinear in each argument,
@@
    are continuous in the associated net convergence structure of the spaces in each case, including multiplication by (fixed) real constants.
is called a bilinear product.
If $X,Y,Z$ are Riesz spaces, we also assume that the bilinear product satisfies:
(iii) It is isotone on the positive cones of $X$ and $Y$, $X^{+}$ and $Y^{+}$ respectively.
(iv) If $\{x_{n}\}_{n\in\mathbb{N}}$ and $\{y_{n}\}_{n\in\mathbb{N}}$ are non-negative sequences in $X$ and $Y$ respectively such that their supremum exists (in $X$ and $Y$ respectively), then:
$\sup_{n\in\mathbb{N}}(x_{n}\bullet y_{i})=(\sup_{n\in\mathbb{N}}x_{n})\bullet y_{i}$
$\sup_{n\in\mathbb{N}}(x_{i}\bullet y_{n})=x_{i}\bullet(\sup_{n\in\mathbb{N}}y_{n}),$
for each fixed $i\in\mathbb{N}$.
For the measure-theoretic structures which the integration theory in the present work will be reduced to, and were first defined by Szabó and Száz (1989) and Szabó and Száz (1988) in a less general setting, we use the next construction:
@@
Let $(\Omega,\mathcal{H})$ be a fixed paved space. A net:
@@
Given these two definitions, we have the main structural object of the present section, defined in analogy to Definition 1.5 of Szabó and Száz (1988):
Definition 2.3 (Integration Structure).
@@
    $(X,\eta_{1}),(Y,\eta_{2}),(Z,\eta_{3})$, vector spaces with (partially ordered, if Riesz spaces) net convergence structures connected by a bilinear product $\bullet$;
will be called an integration structure. If $X,Y,Z$ are Riesz spaces, we shall call the integration structure partially ordered.
With these definitions, we may state the main object of our integration theory:
@@
Let $f:\Omega\rightarrow X$ be a function and $\mu:\mathcal{H}\rightarrow Y$ be an arbitrary set function, both given with respect to an integration structure
@@
Then, if for each $A\in\mathcal{H}$ we have a Fleischer-Szabó-Száz net
$\mathfrak{N}_{A}=\{(\sigma_{\gamma},\tau_{\gamma})\}_{\gamma\in\Gamma_{A}},$
and
$S^{A}_{\gamma}(f,\mu)=\sum_{i\in I_{\gamma}}f(\tau_{\gamma_{i}})\mu(\sigma_{\gamma_{i}}),$
@@
an abstract Riemann sum in $Z$ defined on $\mathfrak{N}_{A}$, such that the following limit exists,
$\int_{A}fd\mu=\lim_{(\Gamma_{A},\eta_{3})}S^{A}_{\gamma}(f,\mu),$
@@
it will be called the net Riemann integral of $f$ with respect to $\mu$ on $A$, or only the net Riemann integral of $f$ if the set function is fixed and the limit exists for every $A\in\mathcal{H}$.
When the context is clear, we shall omit the index $A$ in the Riemann sum $S^{A}_{\gamma}$ for simplicity.
The first relevant comment about this definition is that it is a strict extension of the integral of Szabó and Száz (1988), allowing integration of domains different than the whole domain $\Omega$, incorporating insights from Fleischer (1993), which we also extend, as he imposed further conditions on the integration structure that are not needed in defining the integral.
We also note that this integral is a special case of the following more general construction, inspired by McShane (1969):
Definition 2.5 (Abstract Net Riemann Integral).
Let $U:X\times\mathcal{H}\rightarrow Z$ be a function for $Z$ a vector space equipped with a net convergence structure, and, for each $A\in\mathcal{H}$ and Fleischer-Szabó-Száz net
$\mathfrak{N}_{A}=\{(\sigma_{\gamma},\tau_{\gamma})\}_{\gamma\in\Gamma_{A}},$
put
$S^{U,A}_{\gamma}=\sum_{i\in I_{\gamma}}U(\tau_{\gamma_{i}},\sigma_{\gamma_{i}}),$
an abstract Riemann sum in $Z$ defined on $\mathfrak{N}_{A}$, with respect to $U$. If the following limit exists,
$\int_{A}U=\lim_{(\Gamma_{A},\eta_{3})}S^{U,A}_{\gamma},$
it will be called the abstract net Riemann integral of $U$ on $A$, again omitting $A$ and $U$ in the Riemann sum when it causes no confusion.
We now have the next immediate observation.
Observation 2.1 (Net Riemann Integral from the Abstract Net Riemann Integral).
The net Riemann integral of $f$ with respect to $\mu$ appears as a special case if we pick $U(x,A)=x\mu(A)$, for $A\in\mathcal{A}$, $x$ in the image of $f$ for some $\omega\in\Omega$, and consider the bilinear product of the associated integration structure.
We these constructions in mind, we now prove some general properties of such integrals, focusing on the net Riemann integral. As the first property is quite generic, we prove it for the abstract net Riemann integral. For the next result, we shall need some definitions of convergence structures in product spaces. For that, see O’Brien et al. (2023) and the comments in Hypothesis 1.2. Then, we have, as an immediate consequence of Definition 2.5:
Theorem 2.1.
let $\{(Z_{l},\eta_{l})\}_{l\in L}$ be a collection of net convergence spaces and $Z$ its product equipped with a product convergence structure $\eta$. Then, a function:
$U:X\times\mathcal{H}\rightarrow Z,$
is abstract net Riemann integrable on $A\in\mathcal{H}$ if and only if each $U^{l}:X\times\mathcal{H}\rightarrow Z_{l}$ is net Riemann integrable on $A$ and:
$\int_{A}U=z\iff\int_{A}U^{l}=z_{l},$
where $z\in Z$, $z_{l}=\pi_{l}(z)\in Z_{l}$.
Similarly, from this last result:
Theorem 2.2 (Homomorphic Properties of the abstract net Riemann integral).
Let $\varphi:Z\rightarrow Z^{{}^{\prime}}$ be a continuous homomorphism between the vector spaces $Z,Z^{{}^{\prime}}$ such that $(Z,\eta)$ and $(Z^{{}^{\prime}},\eta^{{}^{\prime}})$ are net convergence spaces. Suppose that $U:X\times\Omega\rightarrow Z$ is abstract net Riemann integrable on $A\in\mathcal{H}$ with respect to a (fixed) net convergence structure. Then, $\varphi\circ U:X\times\Omega\rightarrow Z^{{}^{\prime}}$ is abstract net Riemann integrable on $A$ and
$\int_{A}\varphi\circ U=\varphi(\int_{A}U)$
Proof.
Without loss of generality, we prove this result for $\Omega$ denoting its directed set by $\Gamma$. In this case, Let $S^{U}_{\gamma}$ the corresponding Riemann sum of $U$ for $\gamma\in\Gamma$. As $\varphi$ is a (vector space) homomorphism, we have that:
$S_{\gamma}^{\varphi\circ U}=\sum_{i\in I_{\gamma}}\varphi(U(\tau_{\gamma_{i}},\sigma_{\gamma_{i}}))=\varphi(S_{\gamma}^{U}),$
for each $\gamma\in\Gamma$. As the net $\{S^{U}_{\gamma}\}_{\gamma\in\Gamma}$ has a well defined limit in $Z$ and $\varphi$ is continuous, the right side has a limit in $Z^{{}^{\prime}}$. By the last equality, we then have:
$\int_{\Omega}\varphi(U)=\lim_{(\Gamma,\eta)}S^{\varphi\circ U}_{\gamma}=\lim_{(\Gamma,\eta)}\varphi(S^{U}_{\gamma})=\varphi(\int_{\Omega}U),$
which is the desired result. ∎
From these two general results, and with the choice in the observation above, we get:
Corollary 2.2.1 (Linearity of the net Riemann integral).
If $f_{1},f_{2}$ are net Riemann integrable functions on each $A\in\mathcal{H}$, and with values in a (real) vector space $X$, then $af_{1}+bf_{2}$ is a net Riemann integrable function on each $A\in\mathcal{H}$ for each real constants $a,b\in\mathbb{R}$ and:
$\int_{A}(af_{1}+bf_{2})d\mu=a\int_{A}f_{1}d\mu+b\int_{A}f_{2}d\mu,$
for each $A\in\mathcal{H}$.
@@
To give further properties of the integral in Definition 2.4, we denote by $L_{NR}(\mu)$ the space of all functions $f:\Omega\rightarrow X$ that are net Riemann integrable, on each $A\in\mathcal{H}$, with respect to the (fixed) set function $\mu:\Omega\rightarrow Y$ in the context of a (fixed) integration structure.
Theorem 2.3 (Further Properties of the net Riemann integral).
The net Riemann integral has the following properties compatible with a partially ordered integration structure in Riesz spaces $X,Y,Z$:
@@
    Let $f\in L_{NR}(\mu)$ such that $f\geq 0$ and $\mu\geq 0$. Then,
    $\int_{A}fd\mu\geq 0,$
    for each $A\in\mathcal{H}$, i.e, the integral is a positive map on the space of integrable functions.
@@
    Let $\mu\geq 0$ and $f,g\in L_{NR}(\mu)$ such that:
@@
    $\int_{A}gd\mu\geq\int_{A}fd\mu,$
    for each $A\in\mathcal{H}$, i.e, the integral is an isotone map on $L_{NR}(\mu)$ equipped with the partial ordering inherited from $X$.
  3. (iii)
    Let $\mu\geq 0$ and $f\in L_{NR}(\mu)$ such that $|f|\in L_{NR}(\mu)$, then:
    $|\int_{A}fd\mu|\leq\int_{A}|f|d\mu,$
    for each $A\in\mathcal{H}$.
@@
(i) Let $f\in L_{NR}(\mu)$ such that it, and $\mu$, are non-negative (in $X$ and $Y$ respectively). Let the Riemann sum of $f$ in the fixed integration structure given by $S_{\gamma}(f,\mu)$. By the isotone property of the product and the non-negativity of both $f$ and $\mu$, it follows that:
@@
i.e, the sum is non-negative in $Z$ for every $I_{\gamma}$ and $\gamma\in\Gamma$. Therefore, by taking the limit with respect to the partially ordered net convergence structure $\eta_{3}$, and using that it is locally solid, it follows that: